# G2P: Gaussian-to-Point Attribute Alignment for Boundary-Aware 3D Semantic Segmentation

### v1

@@ 1 Introduction
Point cloud segmentation is a fundamental task that enables comprehensive 3D scene understanding across diverse real-world applications. Despite strong progress [Minkwski, oct, PointTFv3, oneformer3d], the intrinsic sparsity and irregular sampling of point clouds force models to over-rely on coarse geometry, leading to geometric bias that confuses objects with similar shapes but distinct appearances [odin]. For instance, background-adjacent objects (e.g., doors, windows, refrigerators) that are coplanar with walls or floors often become indistinguishable from their surroundings in sparse point clouds (Fig. 1(a)). This geometric bias highlights the need to align appearance cues for accurate discrimination.
Recent advances have improved geometric understanding of point clouds, but existing approaches still face two fundamental limitations. 1) Boundary ambiguity. Boundary-aware approaches [jsenet, tang2022contrastive, bfanet] explicitly model edges, refining object boundaries. However, they remain limited to geometric reasoning and do not take advantage of the appearance information essential to address geometric bias [bundlefusion, 3dpc-survey]. 2) Cross-modal misalignment. 2D-3D fusion approaches [multiview, jaritz2019multiview, chiang2019unified, robert2022learning, vdg-uni3dseg] inject rich image features but suffer from structural mismatches [3dpc-survey]. The discrete nature of point clouds makes wire-like structures hard to distinguish from flat surfaces [mitigating], while projection introduces misalignment and occlusion-induced loss [Kweon2022Joint, odin]. These issues cause segmentation failures even for visually distinct regions, calling for a unified 3D representation that inherently encodes both geometry and appearance.
Among existing representations, 3D Gaussian Splatting (GS) [gaussiansplatting] offers a promising direction that can address both geometric sparsity and appearance deficiency. Unlike point clouds with binary occupancy, Gaussian primitives possess continuous volumetric attributes that encode both geometric structure and appearance properties. Residing in the same 3D coordinate frame as points, these attributes can transfer with less spatial misalignment than 2D-3D fusion [yu2024gaussian, Kweon2022Joint, mitigating, scenesplat]. However, optimization in GS often causes Gaussians to deviate from their initial positions [gaussiansplatting, scaffoldgs, guedon2024sugar, indoorgs], creating a structural mismatch that prevents direct application of Gaussian attributes to point-level segmentation tasks.
@@ 2 Related Works
3D point cloud semantic segmentation has evolved from early per-point methods [pointcnn, pointnet, pointconv] to sophisticated architectures. PointNet++ [pointnetplusplus] introduces hierarchical set abstraction for multi-scale learning, while MinkUNet [Minkwski] transforms point clouds into voxels and employs sparse convolutions to deliver robust and scalable segmentation. Recently, transformer-based architectures such as OctFormer [oct] and Point Transformer v3 [PointTFv3] have demonstrated state-of-the-art performance through efficient self-attention mechanisms.
Beyond supervised learning, self-supervised approaches like PointContrast [pointcontrast] leverage contrastive learning, while Sonata [sonata] demonstrates effective pre-training on large-scale 3D datasets. Despite these advances, purely geometric approaches suffer from boundary ambiguity and geometric bias in sparse point clouds [investigate, bgpseg], particularly where appearance cues become critical for distinguishing geometrically similar objects.
Boundary-Aware Approaches. Starting from Gong et al. [gong2021boundary], recent boundary-aware frameworks such as JSENet [jsenet], CBL [tang2022contrastive], and BFANet [bfanet] explicitly learn boundary features to refine object boundaries in sparse point clouds [edgeaware]. These methods capture geometric discontinuities through dedicated boundary detection modules and edge-aware loss functions. However, they rely mainly on geometric cues and remain limited when segmentation requires appearance cues to address geometric bias arising from similar shapes but different materials.
@@
Cross-Modal 2D-3D Fusion. To address the appearance deficiency, cross-modal 2D-3D fusion approaches leverage rich visual features from RGB images. Multi-view methods such as VMVF [vmvf], 3DMV [3dmv], and MVPNet [jaritz2019multiview] aggregate features from multiple viewpoints through pooling or attention mechanisms. Subsequent studies [bpnet, robert2022learning, bridging, odin] further explore bidirectional projection and pre-trained backbone fusion. However, these approaches are hindered by point-to-pixel alignment errors in occluded regions, projection-induced information loss, and spatial misalignment between sparse 3D distributions and dense 2D grids [Kweon2022Joint, 3dpc-survey]. Consequently, 2D–3D fusion struggles to fully unify appearance and geometry within native 3D space [genova2021learning, interlaced].
@@ 2.2 Gaussian Splatting for 3D Segmentation
With the advent of 3D Gaussian Splatting (GS) [gaussiansplatting], recent works leverage GS for 3D segmentation and scene understanding [gaussiangrouping, flashsplat, segment3dgaussians], and even for open-vocabulary or open-world perception [opengaussian, dcseg, drsplat, refersplat, reasongrounder, identitysplat, panogs]. Beyond improvements in rendering quality, efficiency, and compression, Gaussians encode continuous volumetric geometry and appearance in 3D scenes. By using this potential, numerous methods [feature3dgs, langsplat, trace3d, rethinking] lift CLIP [clip]-/SAM [sam]-derived features or masks into GS and optimize segmentation within the Gaussian representation. These methods focus primarily on view-consistent rendering with segmented Gaussians, which serves a different objective from point cloud segmentation. They also provide no explicit Gaussian-to-point correspondences to preserve the original geometric structure, and instead serve as an intermediate representation for rendering purposes.
UniPre3D [unipre3d] targets point-level tasks, but relies on pixel-wise rendering losses combined with cross-modal fusion to pretrain a backbone network, treating Gaussians as an auxiliary self-supervised signal rather than producing supervised point-level outputs. By contrast, our G2P transfers GS attributes directly to points via explicit Gaussian-to-point alignment under 3D supervision, without requiring recourse to 2D priors or rendering-based losses. Prior evidence shows that these Gaussian attributes encode both appearance and geometric properties [shapesplat, scenesplat, mitigating]. Building on this insight, G2P exploits them as complementary appearance and boundary cues, coupling the appearance richness of Gaussians with the geometric stability of point clouds.
@@ 3.2 Preliminaries
3D GS is a view synthesis method that models a 3D scene as a set of anisotropic Gaussians and renders images through alpha blending. Each Gaussian is parameterized by centroid $\mu^{g}$, opacity $\alpha$, spherical harmonics $SH$, and a covariance matrix $\mathbf{\Sigma}$. The opacity $\alpha$ controls the blending contribution of each Gaussian along the ray, representing its visibility in the rendered view. The spatial shape of a Gaussian is defined by the covariance matrix $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}$ is a diagonal scale matrix. The scale $S$ determines the directional spread and geometric extent of the Gaussian in 3D space. We only use $S$ as a geometric cue and $\alpha$ as appearance information from Gaussian primitives. For additional analysis of Gaussian attributes, please refer to the Supplementary. These choices ensure that G2P focuses on informative attributes that are available for all scenes.
3D GS produces a set of optimized 3D Gaussians whose coordinates are altered during adaptive density optimization. As a result, the original geometric structure of the input point cloud is lost, and point-level semantic labels become unusable [gaussiansplatting, scaffoldgs, guedon2024sugar, indoorgs]. The resulting Gaussian scene exhibits substantially more noise and structural artifacts, such as indistinct edges, when compared to the cleaner input point cloud, making direct use for segmentation unreliable.
@@ 3.3 Gaussian-to-Point Feature Augmentation
To address these fundamental limitations, we propose the Gaussian-to-Point feature augmentation approach. This approach maintains the original point cloud geometry and instead aligns each point with its nearest 3D Gaussians. Each point is then augmented with the associated Gaussian attributes, enabling semantic segmentation while preserving geometric fidelity. To augment the point cloud $\mathcal{P}=\{p_{i}\}_{i=1}^{N}$, where each point $p_{i}=(\mu_{i}^{p},c_{i},n_{i})\in\mathbb{R}^{9}$ consists of coordinates, color, and a normal vector, we transfer attributes from a set of 3D Gaussians $\mathcal{G}=\{g_{j}\}_{j=1}^{M}$. Both sets are first aligned in the same coordinate space. For each point $p_{i}$, we employ a two-stage process to identify its $k$ most relevant Gaussian neighbors. First, to efficiently narrow down candidates, we select all Gaussians whose centroids fall within a Euclidean radius $r$ of $p_{i}$. Second, within this candidate group, we compute the Mahalanobis distance [Mahalanobis] to each Gaussian. Unlike the Euclidean distance that assumes isotropic distributions, it incorporates each Gaussian’s anisotropic shape determined by its scale $S$ and rotation $R$, providing a more physically plausible measure of proximity. The final $k$ neighbors are selected based on the smallest Mahalanobis distances. Accordingly, the proposed weight parameter $w_{ij}$ for each neighbor $g_{j}$ is determined by the inverse of its Mahalanobis distance, normalized across the $k$ neighbors:
@@ 3.5 Gaussian Attribute-guided Learning
To preserve geometric consistency and prevent cross-modal misalignment, our approach takes point clouds as the sole input and receives appearance cues through feature distillation from a teacher encoder. To obtain such appearance representations, we repurpose the Sonata [sonata] architecture to pre-train this teacher encoder without pre-trained weights. While Sonata is a self-supervised framework that independently trains an encoder, in our setup, we configure this framework to learn appearance features. Instead of using the traditional representation $p_{i}=(\mu_{p},c,n)$ with 3D geometric normals $n\in\mathbb{R}^{3}$, we encode each point as $p_{i}=(\mu_{p},c,\alpha)$, where $\alpha\in\mathbb{R}$ denotes opacity, introducing view-consistent appearance cues in place of geometric information. The encoder is trained from scratch on the corresponding dataset (e.g., ScanNet v2 [scannet] or ScanNet++ [scannetpp]), ensuring dataset-specific pre-training for each benchmark. The pre-trained appearance encoder subsequently distills its representations into the 3D segmentation model that operates on point cloud $p_{i}=(\mu_{p},c,n)$.
Our approach incorporates boundary information into semantic segmentation by adopting the boundary-semantic (B-S) block from BFANet [bfanet] and integrating it into the PT v3 [PointTFv3] backbone. The B-S block disentangles the intermediate features into semantic and boundary representations using independent MLPs and enhances semantic features by attending to boundary-aware queries.
@@
VMVF [vmvf] ECCV’20 2D–3D Fusion 76.4
BPNet [bpnet] CVPR’21 69.7
DeepViewAgg [robert2022learning] CVPR’22 71.0
PonderV2 [ponderv2] TPAMI’25 77.0
UniPre3D [unipre3d] CVPR’25 77.6 85.0 92.1
ODIN (Swin-B) [odin] CVPR’24 77.8
MinkUNet [Minkwski] CVPR’19 Geometric 72.3 80.2 90.0
OctFormer [oct] TOG’23 74.3 82.6 90.9
PT v3 [PointTFv3] CVPR’24 77.0 84.3 92.1
BFANet [bfanet] CVPR’25 77.3 84.1 92.2
@@
Semantic supervision loss, $\mathcal{L}_{\text{sem}}$, is applied via a combined loss of cross-entropy and Lovász-softmax [lova], while boundary supervision loss, $\mathcal{L}_{\text{bou}}$, is provided using a combination of binary cross-entropy and Dice loss [diceloss].
@@
MinkUNet [Minkwski] 87.1 96.7 67.0 83.4 92.6 84.5 77.0 81.4 83.7 70.4 65.0 37.8 77.2 64.0 67.2 63.6
OctFormer [oct] 86.1 95.7 70.1 82.7 91.9 83.6 74.1 81.3 83.2 67.7 67.2 34.0 77.4 65.5 65.8 62.9
PT v3 [PointTFv3] 87.7 95.8 73.0 83.2 93.1 83.7 78.9 80.6 84.5 74.6 72.8 41.8 77.9 64.9 68.9 66.8
UniPre3D [unipre3d] 87.5 95.5 72.6 84.4 92.7 84.9 80.0 82.3 85.0 74.6 72.6 38.5 79.0 70.3 76.3 68.6
BFANet [bfanet] 87.2 95.7 72.3 84.7 92.5 83.0 81.2 80.6 84.7 74.0 73.5 38.5 77.5 69.3 70.8 67.3
@@ 4.1 Experiment Setting
Dataset. We evaluate our approach on the ScanNet v2 [scannet, scannet200], ScanNet++ [scannetpp], and Matterport3D [matterport3d] datasets. ScanNet v2, ScanNet200, ScanNet++, and Matterport3D contain 20, 200, 100, and 21 semantic classes, respectively. We report results on the validation splits of all datasets. For 3D Gaussian representations, we adopt the SceneSplat-7K [scenesplat] dataset, where each scene is reconstructed with approximately 1.5M Gaussian primitives. SceneSplat-7K provides GS reconstructions for indoor scenes, but outdoor environments are not available in the dataset.
@@
MinkUNet [Minkwski] CVPR’19 25.0 32.9 80.4
PointContrast [pointcontrast] ECCV’20 26.2
OneFormer3D [oneformer3d] CVPR’24 30.1
PonderV2 [ponderv2] TPAMI’25 32.3 - -
OctFormer [oct] TOG’23 32.6 40.1 83.0
PT v3 [PointTFv3] CVPR’24 34.5 45.4 82.7
UniPre3D [unipre3d] CVPR’25 36.0 46.2 83.7
@@
Evaluation Metric. Following prior works [PointTFv3, bfanet], we adopt three standard metrics for evaluation: mean Intersection over Union (mIoU), mean Accuracy (mAcc), and overall Accuracy (OA), measuring the percentage of correctly classified points across the entire dataset.
Implementation details. Our main training is conducted on a single NVIDIA RTX 3090 GPU for 800 epochs. The appearance encoder pre-training, following Sonata [sonata], is performed separately on an NVIDIA A6000 GPU for 400 epochs per dataset from scratch. We adopt Point Transformer v3 (PT v3) [PointTFv3] as the backbone and incorporate the B-S block from BFANet [bfanet]. The training batch size is set to 4, and we adopt Adam [adamw] as the optimizer with an initial learning rate of 0.003. All other settings follow PT v3, and detailed architectural configurations are reported in the Supplementary. For boundary pseudo-label extraction, we set the scale-based trimming ratio $\tau_{\eta}=0.7$. The radius $r$ for semantic boundary calculation is set to 4 cm. Loss weights are balanced as $\lambda_{d}$ = 0.4, $\lambda_{b}$ = 0.9.

### v2

@@ 1 Introduction
Point cloud segmentation is a fundamental task that enables comprehensive 3D scene understanding across diverse real-world applications. Despite strong progress [7, 48, 53, 26], the intrinsic sparsity and irregular sampling of point clouds force models to over-rely on coarse geometry, leading to geometric bias that confuses objects with similar shapes but distinct appearances [20]. For instance, background-adjacent objects (e.g., doors, windows, refrigerators) that are coplanar with walls or floors often become indistinguishable from their surroundings in sparse point clouds (Fig. 1(a)). This geometric bias highlights the need to align appearance cues for accurate discrimination.
Recent advances have improved geometric understanding of point clouds, but existing approaches still face two fundamental limitations. 1) Boundary ambiguity. Boundary-aware approaches [19, 47, 63] explicitly model edges, refining object boundaries. However, they remain limited to geometric reasoning and do not take advantage of the appearance information essential to address geometric bias [9, 15]. 2) Cross-modal misalignment. 2D-3D fusion approaches [46, 22, 6, 40, 16] inject rich image features but suffer from structural mismatches [15]. The discrete nature of point clouds makes wire-like structures hard to distinguish from flat surfaces [62], while projection introduces misalignment and occlusion-induced loss [28, 20]. These issues cause segmentation failures even for visually distinct regions, calling for a unified 3D representation that inherently encodes both geometry and appearance.
Among existing representations, 3D Gaussian Splatting (GS) [24] offers a promising direction that can address both geometric sparsity and appearance deficiency. Unlike point clouds with binary occupancy, Gaussian primitives possess continuous volumetric attributes that encode both geometric structure and appearance properties. Residing in the same 3D coordinate frame as points, these attributes can transfer with less spatial misalignment than 2D-3D fusion [60, 28, 62, 30]. However, optimization in GS often causes Gaussians to deviate from their initial positions [24, 33, 14, 43], creating a structural mismatch that prevents direct application of Gaussian attributes to point-level segmentation tasks.
@@ 2 Related Works
3D point cloud semantic segmentation has evolved from early per-point methods [29, 36, 51] to sophisticated architectures. PointNet++ [37] introduces hierarchical set abstraction for multi-scale learning, while MinkUNet [7] transforms point clouds into voxels and employs sparse convolutions to deliver robust and scalable segmentation. Recently, transformer-based architectures such as OctFormer [48] and Point Transformer v3 [53] have demonstrated state-of-the-art performance through efficient self-attention mechanisms.
Beyond supervised learning, self-supervised approaches like PointContrast [55] leverage contrastive learning, while Sonata [52] demonstrates effective pre-training on large-scale 3D datasets. Despite these advances, purely geometric approaches suffer from boundary ambiguity and geometric bias in sparse point clouds [56, 11], particularly where appearance cues become critical for distinguishing geometrically similar objects.
Boundary-Aware Approaches. Starting from Gong et al. [13], recent boundary-aware frameworks such as JSENet [19], CBL [47], and BFANet [63] explicitly learn boundary features to refine object boundaries in sparse point clouds [41]. These methods capture geometric discontinuities through dedicated boundary detection modules and edge-aware loss functions. However, they rely mainly on geometric cues and remain limited when segmentation requires appearance cues to address geometric bias arising from similar shapes but different materials.
@@
Cross-Modal 2D-3D Fusion. To address the appearance deficiency, cross-modal 2D-3D fusion approaches leverage rich visual features from RGB images. Multi-view methods such as VMVF [27], 3DMV [10], and MVPNet [22] aggregate features from multiple viewpoints through pooling or attention mechanisms. Subsequent studies [18, 40, 5, 20] further explore bidirectional projection and pre-trained backbone fusion. However, these approaches are hindered by point-to-pixel alignment errors in occluded regions, projection-induced information loss, and spatial misalignment between sparse 3D distributions and dense 2D grids [28, 15]. Consequently, 2D–3D fusion struggles to fully unify appearance and geometry within native 3D space [12, 57].
@@ 2.2 Gaussian Splatting for 3D Segmentation
With the advent of 3D Gaussian Splatting (GS) [24], recent works leverage GS for 3D segmentation and scene understanding [58, 45, 3], and even for open-vocabulary or open-world perception [54, 50, 23, 17, 31, 21, 61]. Beyond improvements in rendering quality, efficiency, and compression, Gaussians encode continuous volumetric geometry and appearance in 3D scenes. By using this potential, numerous methods [65, 38, 44, 67] lift CLIP [39]-/SAM [25]-derived features or masks into GS and optimize segmentation within the Gaussian representation. These methods focus primarily on view-consistent rendering with segmented Gaussians, which serves a different objective from point cloud segmentation. They also provide no explicit Gaussian-to-point correspondences to preserve the original geometric structure, and instead serve as an intermediate representation for rendering purposes.
UniPre3D [49] targets point-level tasks, but relies on pixel-wise rendering losses combined with cross-modal fusion to pretrain a backbone network, treating Gaussians as an auxiliary self-supervised signal rather than producing supervised point-level outputs. By contrast, our G2P transfers GS attributes directly to points via explicit Gaussian-to-point alignment under 3D supervision, without requiring recourse to 2D priors or rendering-based losses. Prior evidence shows that these Gaussian attributes encode both appearance and geometric properties [34, 30, 62]. Building on this insight, G2P exploits them as complementary appearance and boundary cues, coupling the appearance richness of Gaussians with the geometric stability of point clouds.
@@ 3.2 Preliminaries
3D GS is a view synthesis method that models a 3D scene as a set of anisotropic Gaussians and renders images through alpha blending. Each Gaussian is parameterized by centroid $\mu^{g}$, opacity $\alpha$, spherical harmonics $SH$, and a covariance matrix $\mathbf{\Sigma}$. The opacity $\alpha$ controls the blending contribution of each Gaussian along the ray, representing its visibility in the rendered view. The spatial shape of a Gaussian is defined by the covariance matrix $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}$ is a diagonal scale matrix, with the scale denoted as $S$ throughout. The scale $S$ determines the directional spread and geometric extent of the Gaussian in 3D space. We only use $S$ as a geometric cue and $\alpha$ as appearance information from Gaussian primitives. For additional analysis of Gaussian attributes, please refer to the Supplementary. These choices ensure that G2P focuses on informative attributes that are available for all scenes.
3D GS produces a set of optimized 3D Gaussians whose coordinates are altered during adaptive density optimization. As a result, the original geometric structure of the input point cloud is lost, and point-level semantic labels become unusable [24, 33, 14, 43]. The resulting Gaussian scene exhibits substantially more noise and structural artifacts, such as indistinct edges, when compared to the cleaner input point cloud, making direct use for segmentation unreliable.
@@ 3.3 Gaussian-to-Point Feature Augmentation
To address these fundamental limitations, we propose the Gaussian-to-Point feature augmentation approach. This approach maintains the original point cloud geometry and instead aligns each point with its nearest 3D Gaussians. Each point is then augmented with the associated Gaussian attributes, enabling semantic segmentation while preserving geometric fidelity. To augment the point cloud $\mathcal{P}=\{p_{i}\}_{i=1}^{N}$, where each point $p_{i}=(\mu_{i}^{p},c_{i},n_{i})\in\mathbb{R}^{9}$ consists of coordinates, color, and a normal vector, we transfer attributes from a set of 3D Gaussians $\mathcal{G}=\{g_{j}\}_{j=1}^{M}$. Both sets are first aligned in the same coordinate space. For each point $p_{i}$, we employ a two-stage process to identify its $k$ most relevant Gaussian neighbors. First, to efficiently narrow down candidates, we select all Gaussians whose centroids fall within a Euclidean radius $r$ of $p_{i}$. Second, within this candidate group, we compute the Mahalanobis distance [2] to each Gaussian. Unlike the Euclidean distance that assumes isotropic distributions, it incorporates each Gaussian’s anisotropic shape determined by its scale $S$ and rotation $R$, providing a more physically plausible measure of proximity. The final $k$ neighbors are selected based on the smallest Mahalanobis distances. Accordingly, the proposed weight parameter $w_{ij}$ for each neighbor $g_{j}$ is determined by the inverse of its Mahalanobis distance, normalized across the $k$ neighbors:
@@ 3.5 Gaussian Attribute-guided Learning
To preserve geometric consistency and prevent cross-modal misalignment, our approach takes point clouds as the sole input and receives appearance cues through feature distillation from a teacher encoder. To obtain such appearance representations, we repurpose the Sonata [52] architecture to pre-train this teacher encoder without pre-trained weights. While Sonata is a self-supervised framework that independently trains an encoder, in our setup, we configure this framework to learn appearance features. Instead of using the traditional representation $p_{i}=(\mu_{p},c,n)$ with 3D geometric normals $n\in\mathbb{R}^{3}$, we encode each point as $p_{i}=(\mu_{p},c,\alpha)$, where $\alpha\in\mathbb{R}$ denotes opacity, introducing view-consistent appearance cues in place of geometric information. The encoder is trained from scratch on the corresponding dataset (e.g., ScanNet v2 [8] or ScanNet++ [59]), ensuring dataset-specific pre-training for each benchmark. The pre-trained appearance encoder subsequently distills its representations into the 3D segmentation model that operates on point cloud $p_{i}=(\mu_{p},c,n)$.
Our approach incorporates boundary information into semantic segmentation by adopting the boundary-semantic (B-S) block from BFANet [63] and integrating it into the PT v3 [53] backbone. The B-S block disentangles the intermediate features into semantic and boundary representations using independent MLPs and enhances semantic features by attending to boundary-aware queries.
@@
VMVF [27] ECCV’20 2D–3D Fusion 76.4
BPNet [18] CVPR’21 69.7
DeepViewAgg [40] CVPR’22 71.0
PonderV2 [66] TPAMI’25 77.0
UniPre3D [49] CVPR’25 77.6 85.0 92.1
ODIN (Swin-B) [20] CVPR’24 77.8
MinkUNet [7] CVPR’19 Geometric 72.3 80.2 90.0
OctFormer [48] TOG’23 74.3 82.6 90.9
PT v3 [53] CVPR’24 77.0 84.3 92.1
BFANet [63] CVPR’25 77.3 84.1 92.2
@@
Semantic supervision loss, $\mathcal{L}_{\text{sem}}$, is applied via a combined loss of cross-entropy and Lovász-softmax [1], while boundary supervision loss, $\mathcal{L}_{\text{bou}}$, is provided using a combination of binary cross-entropy and Dice loss [35].
@@
MinkUNet [7] 87.1 96.7 67.0 83.4 92.6 84.5 77.0 81.4 83.7 70.4 65.0 37.8 77.2 64.0 67.2 63.6
OctFormer [48] 86.1 95.7 70.1 82.7 91.9 83.6 74.1 81.3 83.2 67.7 67.2 34.0 77.4 65.5 65.8 62.9
PT v3 [53] 87.7 95.8 73.0 83.2 93.1 83.7 78.9 80.6 84.5 74.6 72.8 41.8 77.9 64.9 68.9 66.8
UniPre3D [49] 87.5 95.5 72.6 84.4 92.7 84.9 80.0 82.3 85.0 74.6 72.6 38.5 79.0 70.3 76.3 68.6
BFANet [63] 87.2 95.7 72.3 84.7 92.5 83.0 81.2 80.6 84.7 74.0 73.5 38.5 77.5 69.3 70.8 67.3
@@ 4.1 Experiment Setting
Dataset. We evaluate our approach on the ScanNet v2 [8, 42], ScanNet++ [59], and Matterport3D [4] datasets. ScanNet v2, ScanNet200, ScanNet++, and Matterport3D contain 20, 200, 100, and 21 semantic classes, respectively. We report results on the validation splits of all datasets. For 3D Gaussian representations, we adopt the SceneSplat-7K [30] dataset, where each scene is reconstructed with approximately 1.5M Gaussian primitives. SceneSplat-7K provides GS reconstructions for indoor scenes, but outdoor environments are not available in the dataset.
@@
MinkUNet [7] CVPR’19 25.0 32.9 80.4
PointContrast [55] ECCV’20 26.2
OneFormer3D [26] CVPR’24 30.1
PonderV2 [66] TPAMI’25 32.3 - -
OctFormer [48] TOG’23 32.6 40.1 83.0
PT v3 [53] CVPR’24 34.5 45.4 82.7
UniPre3D [49] CVPR’25 36.0 46.2 83.7
@@
Evaluation Metric. Following prior works [53, 63], we adopt three standard metrics for evaluation: mean Intersection over Union (mIoU), mean Accuracy (mAcc), and overall Accuracy (OA), measuring the percentage of correctly classified points across the entire dataset.
Implementation details. Our main training is conducted on a single NVIDIA RTX 3090 GPU for 800 epochs. The appearance encoder pre-training, following Sonata [52], is performed separately on an NVIDIA A6000 GPU for 400 epochs per dataset from scratch. We adopt Point Transformer v3 (PT v3) [53] as the backbone and incorporate the B-S block from BFANet [63]. The training batch size is set to 4, and we adopt Adam [32] as the optimizer with an initial learning rate of 0.003. All other settings follow PT v3, and detailed architectural configurations are reported in the Supplementary. For boundary pseudo-label extraction, we set the scale-based trimming ratio $\tau_{\eta}=0.7$. The radius $r$ for semantic boundary calculation is set to 4 cm. Loss weights are balanced as $\lambda_{d}$ = 0.4, $\lambda_{b}$ = 0.9.